At Midland Company's break-even point of 9,400 units, fixed costs are $225,600 and variable costs are $601,600 in total. The unit sales price is:

$112.

$24.

$88.

$40.

$64.
Answer #1

Variable cost per unit=(601600/9400)=$64

Breakeven=Fixed expenses/Contribution margin

Contribution margin=225600/9400

=$24

Contribution margin=Sales-Variable cost  

Hence sales price=24+64

=$88